# Simultaneous Ground Reaction Force and State Estimation via Constrained Moving Horizon Estimation

### v1

@@ Abstract
Accurate ground reaction force (GRF) estimation can significantly improve the adaptability of legged robots in various real-world applications. For instance, with estimated GRF and contact kinematics, the locomotion control and planning assist the robot in overcoming uncertain terrains. The canonical momentum-based methods, formulated as nonlinear observers, do not fully address the noisy measurements and the dependence between floating base states and the generalized momentum dynamics. In this paper, we present a simultaneous ground reaction force and state estimation framework for legged robots, which systematically addresses the sensor noise and the coupling between states and dynamics. With the floating base orientation estimated separately, a decentralized Moving Horizon Estimation (MHE) method is implemented to fuse the robot dynamics, proprioceptive sensors, exteroceptive sensors, and deterministic contact complementarity constraints in a convex windowed optimization. The proposed method is shown to be capable of providing accurate GRF and state estimation on several legged robots, including the open-source educational planar bipedal robot STRIDE and quadrupedal robot Unitree Go1, with a frequency of 200Hz and a past time window of 0.04s.
@@ I Introduction
Installing contact sensors, such as pressure sensors [4], on legged robots is a common and straightforward method for providing a binary indicator of contact states, however, only a few robots are equipped with multi-axis force/torque (F/T) sensors [5, 6], due to their high cost, weight, and fragility. GRF estimation using a general sensor setup, i.e., IMUs, joint encoders, joint current sensors, stereo cameras, and foot contact sensors, can offer a more practical, efficient, and effective solution.
@@
However, unlike manipulators, the dynamics of legged robots depend on floating base states - a relationship that has not been fully addressed in previous works. In [8], [9], and [10], GRF and floating base states are estimated separately, leading to a disconnection that limits the ability to fully capture the system’s dynamics. Furthermore, in scenarios involving rigid contact, contact complementarity constraints [12] have been shown to enhance estimation accuracy during sliding and collision events [13]. Therefore, an optimization-based estimator that simultaneously estimates GRF and states, while incorporating constraint information, offers an ideal solution.
In this study, we extend our previous decentralized MHE framework in [14] to achieve simultaneous estimation of the GRF and floating-base states of the legged robot, and we demonstrate its superior performance compared to existing methods. The framework structure is illustrated in Fig. 1, with orientation estimated separately by another orientational filer or an Attitude and Heading Reference System (AHRS), the MHE is formulated and solved as a Quadratic Program (QP), which allows real-time estimation. This study makes the following contributions:
  • A formulation of a novel state estimation framework using decentralized MHE [14], which achieves simultaneous ground reaction force and state estimation, fusing proprioceptive, exteroceptive sensors and physical constraints in real-time.
  • Evaluation of the proposed method on the open-sourced planar bipedal robot STRIDE [15] and the commercial quadrupedal robot Unitree Go1 [4].
  • Comparisons of the proposed method with conventional approaches, highlighting improved performance due to the integration of exteroceptive sensors and contact constraints.
@@ II Related Work
To address the challenges associated with sensor design, sensorless methods have tackled the GRF estimation problem using a dynamics-based approach. In [18], the disturbance caused by GRFs on the humanoid robot HRP-4 is directly obtained by comparing the full-body dynamics with measured or estimated accelerations of the generalized coordinate. The effectiveness of this method is highly dependent on the accuracy of the sensor data. Momentum-based methods mitigate this requirement by estimating the GRF using momentum dynamics. [8], [9] estimated the GRF on the quadrupedal robot with a Momentum Based Observer (MBO). However, these methods handle noise similarly to low-pass filters, without accurately addressing the sensor noise or bias. Additionally, despite that the dynamics and configuration of the robot depend on the floating-base states, these methods estimate the floating-base states and GRF separately using a loosely coupled approach, which is both redundant and sub-optimal.
Considering modeling uncertainties and sensor noise, [11] proposed a momentum-based Disturbance Kalman Filter (DKF) for force estimation on a robotic manipulator. To further address the correlation between floating-base estimation and force estimation, [5] developed a constrained Quadratic Program (QP) based estimator that leverages both full-body dynamics and 3-axis F/T sensors on Atlas to estimate contact forces and robot states. However, the estimator performance is only partially demonstrated and perhaps limited by the capabilities of the F/T sensor, since the results focus solely on GRF estimation in the vertical direction.
Recent research in [19] formulated a nonlinear MHE to achieve a simultaneous estimation of forces, parameters, and states using nonholonomic-constrained full-body dynamics [20]. However, due to the computational demands of nonlinear optimization, the algorithm operates offline, and no results of the estimated GRFs were provided. In contrast, our approach aims to achieve real-time simultaneous GRF and state estimation using Decentralized MHE framework [14]. This method integrates momentum dynamics, exteroceptive sensors, and contact complementarity constraints to enhance estimation performance.
@@ III-A Moving Horizon Estimation Formulation
A general description of the dynamic system involves a nonlinear difference equation of its dynamics:
$x^{+}=f(x,u)+\delta^{x},$ (1)
and a measurement equation:
$y=h(x)+\delta^{y},$ (2)
where $(\cdot)^{+}$ denotes the updated state after a fixed interval $\Delta t$, $f(\cdot)$ and $h(\cdot)$ denote the process model and measurement model, respectively. $x$, $u$, and $y$ denote the state, control input, and measurement of the system, respectively. $\delta^{x}\sim\mathcal{N}(0,Q)$ and $\delta^{y}\sim\mathcal{N}(0,R)$ represent the process noise and measurement noise subject to zero-mean Gaussian distributions. Additionally, the legged robot is often subject to various physical constraints, which can be generally described as state constraints:
$z(x)\leq 0.$ (3)
The MHE solves a Maximum A Posteriori (MAP) problem for a trajectory of states $x_{[T-N:T]}$. It incorporates the control inputs $u_{[T-N:T-1]}$ and measurements $y_{[T-N,T]}$ within a receding window of size $N+1$, spanning from time index $T-N$ to $T$. The notation $(\cdot)_{[i,j]}$ denotes the sequence of vectors starting from the time index $i$ and ending at the time index $j$. The solution at each time step is obtained by solving the MAP problem, subject to the physical constraints in (3):
$\underset{x_{[T-N,T]}}{\text{argmax}}$ $\ {P(x_{[T-N,T]}|u_{[T-N,T-1]},y_{[T-N,T]})},$ (4)
s.t. $\quad x_{[T-N,T]}\in\mathcal{X},$
where $P$ denotes the posterior probability, and $\mathcal{X}$ represents the feasible state region considering the imposed constraints. The MHE, formulated as an optimization problem, minimizes the minus log-likelihood of (4) to compute the optimal estimation of $\mathbf{X}_{[T-N:T]}$:
$\mathbf{X}_{[T-N:T]}=\begin{bmatrix}x_{[T-N,T]}&\delta^{x}_{[T-N,T-1]}&\delta^{y}_{[T-N,T]}\end{bmatrix},$ (5)
considering the process model (1), measurement model (2) and physical constraints (3) into this optimization problem:
$\min_{\mathbf{X}_{[T-N,T]}}\ \Gamma(x_{T-N})$
$+\sum_{k=T-N}^{T-1}||\delta^{x}_{k}||_{Q_{k}}+\sum_{k=T-N}^{T}||\delta^{y}_{k}||_{R_{k}},$
@@
where $\Gamma(x_{T-N})$ denotes the arrival cost. This term enables the processing of a finite horizon of measurements by incorporating all information prior to the current horizon into the arrival cost. The standard approach computes the arrival cost using an Extended Kalman Filter [21], and recent research also proposed a probabilistic approach to derive this term from the Hessian matrix of (IV-E) [22]. In this paper, we will adopt the arrival cost calculation method from our previous work [14], which involves solving the Schur complement of the KKT matrix for the equality-constrained MHE.
@@ III-B Orientation Estimation Formulation
Following prior practices in [23], [24], the decentralized MHE [14] utilizes a standalone and lightweight estimator for orientation estimation. Recently, off-the-shelf Attitude and Heading Reference Systems (AHRS) have demonstrated their ability to provide reliable orientation measurements $\hat{R}_{\text{WB}}$, relative to gravity and the Earth’s magnetic field [25]. For legged robots in static contact scenarios, filter-based state estimation methods, such as the Invariant-Extended Kalman Filter (InEKF) [26], incorporate additional sensor data to deliver accurate orientation estimates. In our previous work [14], a vision-aided, quaternion-based Iterated Kalman Filter (QEKF) [27] was developed to provide a general robotics solution for orientation estimation. In this work, we retain the orientation estimator that we developed in [14]. Generally, without a high-quality IMU sensor, vision information is the primary additional information needed to provide accurate orientation estimation.
@@ III-C Generalized Momentum Dynamics
The configuration of a legged robot is typically described using the generalized coordinates $q\in SE(3)\times\mathbb{R}^{n}$ that include the floating base states:
@@
where $p$, $[\phi\ \theta\ \psi]$, and $\alpha\in\mathbb{R}^{n}$ represent the floating base position, Euler angles in the world frame, and joint angles, respectively. $n$ denotes the dimension of the joint states. Consider the full-body floating-base dynamics of a legged robot:
$M(q)\ddot{q}+C(q,\dot{q})\dot{q}+G(q)=Bu+\sum_{i}J_{i}(q)^{T}f_{i},$ (10)
where $M(q)\in\mathbb{R}^{(6+n)\times(6+n)}$, $C(q,\dot{q})\in\mathbb{R}^{(6+n)\times(6+n)}$ and $G(q)\in\mathbb{R}^{(6+n)}$ represent inertial matrix, Coriolis matrix, and gravitational vector, respectively. $B$ denotes a constant matrix that maps the joint torque vector $u$ to the generalized force vector. $\dot{q}$ and $\ddot{q}$ denote the first and second-time derivatives of the generalized coordinates, and $f_{i}$ and $J_{i}(q)$ denote the contact GRF and its associated coordinate Jacobian, respectively. Specifically, for legged robots with point feet, the GRF $f_{i}$ acting on each individual foot $i$ is considered to have three degrees of freedom (DOF).
Given the torques applied at the actuators $u$ and generalized coordinates accelerations $\ddot{q}$, the expected trajectory outcomes based on the dynamics model can be compared to the actual results, with any discrepancies attributed to external wrenches. The momentum-based method leverages the properties of the Coriolis matrix, eliminating the need for noisy estimates or measurements of the second-order derivative $\ddot{q}$. The generalized momentum and its derivatives are defined respectively as:
$m$ $=M(q)\dot{q},$ (11)
$\dot{m}$ $=\dot{M(q)}\dot{q}+M(q)\ddot{q}.$ (12)
@@ IV Simultaneous Force and State Estimation
This section introduces the proposed method for estimating GRF alongside the floating-base states of the robot. We formulate a Decentralized Moving Horizon Estimation based on [14], and incorporate additional momentum and contact constraint. The state (5), process model (7), measurement model (6), and the physical constraints (8) in (IV-E) are detailed.
@@ IV-A Decentralized Estimation
For general estimation tasks, a set of robot-centric states is selected to describe the motion of the legged robot, enabling simultaneous estimation of the robot’s base position, velocity, and orientation. In the decentralized framework used in [23], [24], the base orientation is estimated separately from the position and velocity, using a standalone estimator. Considering the unreliable performance of magnetometers in robotic applications and the terrain dependency of kinematics-based measurements, we proposed a vision-aided iterated QEKF in our previous work [14]. This method iteratively fuses IMU readings with delayed VO corrections through synchronization and trajectory updates, consistently providing accurate orientation estimates. The estimation result $\hat{R}_{\text{WB}}$ is used to simplify the following process and measurement models to linear time-varying (LTV) constraints of the MHE, as illustrated in Fig. 1. This decentralization transforms the MHE into a convex QP problem. The notation $\hat{(\cdot)}$ and $\tilde{(\cdot)}$ represent the estimated and measured quantity, respectively.
@@ IV-B Process Model
where $p$, $v$, and $f$ denote the base position, base velocity, and GRF, respectively, in the world frame denoted by $W$. $b_{a}$ denotes the accelerometer bias in the body frame $B$. $m$ denotes the generalized momentum. For simplicity, only one of the $N$ legs is included in the notation. The actual number of feet considered depends on the specific robot and its locomotion behavior.
@@
$x^{+}=\begin{bmatrix}p+v\Delta t+\frac{1}{2}\big{(}\hat{R}_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t^{2}\\ v+\big{(}\hat{R}_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t\\ b_{a}\\ \textstyle m+\Delta t(\dot{m})\\ f_{i}\end{bmatrix}+\delta_{x},$ (16)
@@
$\dot{m}=C_{1}^{\intercal}v+C_{2}^{\intercal}\begin{bmatrix}\tilde{\omega}^{\intercal}&\tilde{\dot{\alpha}}^{\intercal}\end{bmatrix}^{\intercal}-G+B\tilde{u}+\sum_{i}{J_{i}^{\intercal}f_{i}},$ (17)
where $C_{1}\in\mathbb{R}^{3\times(n+6)}$ and $C_{2}\in\mathbb{R}^{(n+3)\times(n+6)}$ are sub-blocks of the Coriolis Matrix $C(q,\dot{q})$, evaluated at the estimated generalized coordinates at the associated time:
$C=\begin{bmatrix}C_{1}^{\intercal}&C_{2}^{\intercal}\end{bmatrix}^{\intercal}.$ (18)
Considering the relatively small magnitude of the Coriolis matrix, the effects of linearization can be mitigated and treated as model uncertainties.

### v2

@@ Abstract
Accurate ground reaction force (GRF) estimation can significantly improve the adaptability of legged robots in various real-world applications. For instance, with estimated GRF and contact kinematics, the locomotion control and planning assist the robot in overcoming uncertain terrains. The canonical momentum-based methods, formulated as nonlinear observers, do not fully address the noisy measurements and the dependence between floating-base states and the generalized momentum dynamics. In this paper, we present a simultaneous ground reaction force and state estimation framework for legged robots, which systematically addresses the sensor noise and the coupling between states and dynamics. With the floating base orientation estimated separately, a decentralized Moving Horizon Estimation (MHE) method is implemented to fuse the robot dynamics, proprioceptive sensors, exteroceptive sensors, and deterministic contact complementarity constraints in a convex windowed optimization. The proposed method is shown to be capable of providing accurate GRF and state estimation on several legged robots, including the custom-designed humanoid robot Bucky, the open-source educational planar bipedal robot STRIDE, and the quadrupedal robot Unitree Go1, with a frequency of 200Hz and a past time window of 0.04s.
@@ I Introduction
Installing contact sensors, such as pressure sensors [4], on legged robots is a common and straightforward method for providing a binary indicator of contact states. However, only a few robots are equipped with multi-axis force/torque (F/T) sensors [5, 6], due to their high cost, weight, and fragility. GRF estimation using a general sensor setup, i.e., IMUs, joint encoders, joint current sensors, stereo cameras, and foot contact sensors, can offer a more practical, efficient, and effective solution.
@@
However, unlike manipulators, legged robots have dynamics that depend on floating base states - a relationship that previous works have not fully addressed. [8], [9], and [10] estimated GRF and floating-base states separately, creating a disconnect that limits the ability to fully capture the system’s dynamics. Furthermore, in scenarios involving rigid contact, contact complementarity constraints [12] have been shown to enhance estimation accuracy during sliding and collision events [13]. Therefore, an optimization-based estimator that simultaneously estimates GRF and states, while incorporating constraint information, offers an ideal solution.
Figure 1: Illustration of the proposed decentralized estimation framework for estimation on various legged robots, including the humanoid robot Bucky (left) and the quadrupedal robot B1 (right). The experiment video is available at: https://youtu.be/Bih7cslSkTo.
In this study, we extend our previous decentralized MHE framework in [14] to achieve simultaneous estimation of the GRF and floating-base states of the legged robot, and we demonstrate its superior performance compared to existing methods. The framework structure is illustrated in Fig. 1. With orientation estimated separately by another orientational filer or an Attitude and Heading Reference System (AHRS), the MHE is formulated and solved as a Quadratic Program (QP), which allows real-time estimation. This study makes the following contributions:
  • A novel state estimation framework based on decentralized MHE [14], enabling simultaneous GRF and state estimation by fusing proprioceptive and exteroceptive sensors with physical constraints in real time.
  • Evaluation of the proposed method on the custom-designed humanoid robot Bucky, the open-sourced planar bipedal robot STRIDE [15] and the commercial quadrupedal robot Unitree Go1 [4].
  • Comparison of the proposed method with conventional approaches, highlighting performance improvements from exteroceptive sensors and contact constraints.
@@ II Related Work
To address the challenges associated with sensor design, sensorless methods have tackled the GRF estimation problem using a dynamics-based approach. In [18], the disturbance caused by GRFs on the humanoid robot HRP-4 is directly obtained by comparing the full-body dynamics with measured or estimated accelerations of the generalized coordinate. The effectiveness of this method is highly dependent on the accuracy of the sensor data. Momentum-based methods mitigate this requirement by estimating the GRF using momentum dynamics. [8], [9] estimated the GRF on the quadrupedal robot with a Momentum-Based Observer (MBO). However, these methods handle noise similarly to low-pass filters, without accurately addressing the sensor noise or bias. Additionally, despite that the dynamics and configuration of the robot depend on the floating-base states, these methods estimate the floating-base states and GRF separately using a loosely coupled approach, which is both redundant and sub-optimal.
Considering modeling uncertainties and sensor noise, [11] proposed a momentum-based Disturbance Kalman Filter (DKF) for force estimation on a robotic manipulator. To further address the correlation between floating-base estimation and force estimation, [5] developed a constrained Quadratic Program (QP) based estimator that leverages both full-body dynamics and 3-axis F/T sensors on Atlas to estimate contact forces and robot states. However, the estimator performance is only partially demonstrated and potentially limited by the capabilities of the F/T sensor, since the results focus solely on GRF estimation in the vertical direction.
Recent research in [19] formulated a nonlinear MHE to achieve a simultaneous estimation of forces, parameters, and states using holonomic-constrained full-body dynamics [20]. However, due to the computational demands of nonlinear optimization, the algorithm runs offline, and no estimated GRF results are provided. In contrast, our approach aims to achieve real-time simultaneous GRF and state estimation using decentralized MHE [14]. This method integrates momentum dynamics, exteroceptive sensors, and contact complementarity constraints to enhance estimation performance.
@@ III-A Moving Horizon Estimation Formulation
A dynamic system is generally described by a nonlinear difference equation and a measurement equation:
$x^{+}=f(x,u)+\delta^{x},\quad y=h(x)+\delta^{y},$ (1)
where $(\cdot)^{+}$ denotes the updated state after a fixed interval $\Delta t$, $f(\cdot)$ and $h(\cdot)$ denote the process model and measurement model, respectively. $x$, $u$, and $y$ denote the state, control input, and measurement of the system, respectively. $\delta^{x}\sim\mathcal{N}(0,Q)$ and $\delta^{y}\sim\mathcal{N}(0,R)$ represent the process noise and measurement noise subject to zero-mean Gaussian distributions. Additionally, legged robots are often subject to various physical constraints, which can generally be expressed as state constraints $z(x)\leq 0$.
The MHE solves a Maximum A Posteriori (MAP) problem for a trajectory of states $x_{[T-N:T]}$. It incorporates the control inputs $u_{[T-N:T-1]}$ and measurements $y_{[T-N,T]}$ within a receding window of size $N+1$, spanning from time index $T-N$ to $T$. The notation $(\cdot)_{[i,j]}$ denotes the sequence of vectors starting from the time index $i$ and ending at the time index $j$. The solution at each time step is obtained by solving the MAP problem, subjects to the physical constraints:
$\underset{x_{[T-N,T]}\in\mathcal{X}}{\text{argmax}}$ $\ {P(x_{[T-N,T]}|u_{[T-N,T-1]},y_{[T-N,T]})}\textstyle,$ (2)
where $P$ denotes the posterior probability, and $\mathcal{X}$ represents the feasible state region considering the imposed constraints. The MHE, formulated as an optimization problem, minimizes the negative log-likelihood of (2) to compute the optimal estimation of $\mathbf{X}_{[T-N:T]}$:
$\mathbf{X}_{[T-N:T]}=\begin{bmatrix}x_{[T-N,T]}&\delta^{x}_{[T-N,T-1]}&\delta^{y}_{[T-N,T]}\end{bmatrix},\textstyle$ (3)
integrating the process and measurement model along with the physical constraints into the optimization:
$\min_{\mathbf{X}_{[T-N,T]}}\ \Gamma(x_{T-N})\textstyle$
$+\textstyle{\sum}_{k=T-N}^{T-1}||\delta^{x}_{k}||_{Q_{k}}+\textstyle{\sum}_{k=T-N}^{T}||\delta^{y}_{k}||_{R_{k}},\textstyle$
@@
where $\Gamma(x_{T-N})$ denotes the arrival cost. This term enables the processing of a finite horizon of measurements by incorporating all information prior to the current horizon into the arrival cost. The standard approach computes the arrival cost using an Extended Kalman Filter [21], and recent research also proposes a probabilistic approach to derive this term from the Hessian matrix of (III-A) [22]. In this paper, we adopt the arrival cost calculation method from our previous work [14], which involves solving the Schur complement of the KKT matrix for the equality-constrained MHE.
@@ III-B Orientation Estimation Formulation
Following prior practices in [23], [24], the decentralized MHE [14] utilizes a standalone and lightweight estimator for orientation estimation. Recently, off-the-shelf Attitude and Heading Reference Systems (AHRS) have demonstrated their ability to provide reliable orientation measurements $\hat{R}_{\mathcal{WB}}$, relative to gravity and the Earth’s magnetic field [25]. For legged robots in static contact scenarios, filter-based state estimation methods, such as the Invariant-Extended Kalman Filter (InEKF) [26], incorporate additional sensor data to deliver accurate orientation estimates. Generally, without a high-quality IMU sensor, vision information is the primary additional information needed to provide accurate orientation estimation. In our previous work [14], a vision-aided, quaternion-based Iterated Kalman Filter (QEKF) [27] was developed to provide a general robotics solution for orientation estimation. In this work, we retain the orientation estimator that we developed in [14].
@@ III-C Generalized Momentum Dynamics
The configuration of a legged robot is typically described using the generalized coordinates $q\in SE(3)\times\mathbb{R}^{n}$ that include the floating-base states:
@@
where $p$, $[\phi\ \theta\ \psi]$, and $\alpha\in\mathbb{R}^{n}$ represent the floating-base position, Euler angles in the world frame, and joint angles, respectively. $n$ denotes the dimension of the joint states. Consider the full-body dynamics of a legged robot:
$M(q)\ddot{q}+C(q,\dot{q})\dot{q}+G(q)=Bu+\textstyle{\sum}_{i}J_{i}(q)^{T}f_{i},\textstyle$ (8)
where $M(q)\in\mathbb{R}^{(6+n)\times(6+n)}$, $C(q,\dot{q})\in\mathbb{R}^{(6+n)\times(6+n)}$ and $G(q)\in\mathbb{R}^{(6+n)}$ represent inertia matrix, Coriolis matrix, and gravitational vector, respectively. $B$ denotes a constant matrix that maps the joint torque vector $u$ to the generalized force vector. $\dot{q}$ and $\ddot{q}$ denote the first and second-time derivatives of the generalized coordinates, and $f_{i}$ and $J_{i}(q)$ denote the contact GRF and its associated coordinate Jacobian, respectively. Specifically, for legged robots with point feet, the GRF $f_{i}$ acting on each individual foot $i$ is considered to have three degrees of freedom (DOF).
Given the torques applied at the actuators $u$ and generalized coordinates accelerations $\ddot{q}$, the expected trajectory outcomes based on the dynamics model can be compared to the actual results, with any discrepancies attributed to external wrenches. The momentum-based method leverages the properties of the Coriolis matrix, eliminating the need for noisy estimates or measurements of the second-order derivative $\ddot{q}$. The generalized momentum $m$ and its derivatives $\dot{m}$ are defined respectively as:
$m=M(q)\dot{q},\quad\dot{m}=\dot{M(q)}\dot{q}+M(q)\ddot{q}.$ (9)
@@ IV Simultaneous Force and State Estimation
This section introduces the proposed method for estimating GRF alongside the floating-base states of the robot. We formulate a decentralized Moving Horizon Estimation [14], incorporating additional momentum and contact constraints. The state (3), process model (5), measurement model (4), and the physical constraints (6) in (III-A) are detailed.
@@ IV-A Decentralized Estimation
For general estimation tasks, a set of robot-centric states are selected to describe the motion of the legged robot, enabling simultaneous estimation of the robot’s base position, velocity, and orientation. In the decentralized framework used in [23], [24], the base orientation is estimated separately from the position and velocity, using a standalone estimator. Considering the unreliable performance of magnetometers in robotic applications and the terrain dependency of kinematics-based measurements, we proposed a vision-aided iterated QEKF in our previous work [14]. This method iteratively fuses IMU readings with delayed VO corrections through synchronization and trajectory updates, consistently providing accurate orientation estimates. The estimation result $\hat{R}_{\mathcal{WB}}$ is used to simplify the following process and measurement models to linear time-varying (LTV) constraints of the MHE, as illustrated in Fig. 1. This decentralization transforms the MHE into a convex QP problem. The notation $\hat{(\cdot)}$ and $\tilde{(\cdot)}$ represent the estimated and measured quantity, respectively.
@@ IV-B Process Model
where $p$, $v$, and $f$ denote the base position, base velocity, and GRF, respectively, in the world frame denoted by $\mathcal{W}$. $b_{a}$ denotes the accelerometer bias in the body frame $\mathcal{B}$. $m$ denotes the generalized momentum. For simplicity, only one of the $N$ legs is included in the notation. The actual number of feet considered depends on the specific robot and its locomotion behavior.
@@
$x^{+}=\begin{bmatrix}p+v\Delta t+\frac{1}{2}\big{(}\hat{R}_{\mathcal{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t^{2}\\ v+\big{(}\hat{R}_{\mathcal{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t\\ b_{a}\\ \textstyle m+\Delta t\dot{m}\\ f_{i}\end{bmatrix}+\delta_{x},$ (13)
@@
$\dot{m}=C_{1}^{\intercal}v+C_{2}^{\intercal}\begin{bmatrix}\tilde{\omega}^{\intercal}&\tilde{\dot{\alpha}}^{\intercal}\end{bmatrix}^{\intercal}-G+B\tilde{u}+\textstyle{\sum}_{i}{J_{i}^{\intercal}f_{i}},$ (14)
where $C_{1}\in\mathbb{R}^{3\times(n+6)}$ and $C_{2}\in\mathbb{R}^{(n+3)\times(n+6)}$ are sub-blocks of the Coriolis matrix $C=\begin{bmatrix}C_{1}^{\intercal}&C_{2}^{\intercal}\end{bmatrix}^{\intercal}$, evaluated at the estimated generalized coordinates for the corresponding time. Considering the relatively small magnitude of the Coriolis matrix, the effects of linearization can be mitigated and treated as model uncertainties.